# Deep Smart Contract Intent Detection

### v2

@@ V-C Multi-label Classification
Formulas 13 to 18 illustrate the computation process after a row vector $\bm{f_{i}}$ is input into an LSTM cell, where $i\in\mathbb{N}^{+}$ and $i\leq p$. Here, $\Theta$ denotes the cell state vector, and $\widetilde{\Theta}$ signifies the cell input activation vector. The initial state $\Theta_{0}$ is initialized as a zero vector with a dimensionality of $h$ (hidden units per LSTM cell). The various gates within the LSTM cell are represented as follows: $\mathrm{G}^{\mathsf{f}}_{i}$ is the forget gate, $\mathrm{G}^{\mathsf{u}}_{i}$ is the update (or input) gate, and $\mathrm{G}^{\mathsf{o}}_{i}$ is the output gate. The weight matrices $\bm{W}\in\mathbb{R}^{h\times m}$ and $\bm{U}\in\mathbb{R}^{h\times h}$, along with the bias vector $\bm{b}\in\mathbb{R}^{h}$, are parameters learned during the training phase. The activation functions denoted by $\alpha$, $\beta$, and $\gamma$ can be $\mathrm{tanh}$ or $\mathrm{sigmoid}$. The hidden state vector $\bm{h_{i}}$, which is also the output of the LSTM cell, has a dimensionality of $h$, with the initial hidden state $\bm{h_{0}}$ being initialized as a zero vector. The operator $\odot$ denotes the Hadamard product (element-wise product) [49].
@@ VI-B Evaluation Metrics
Using the values of TP, TN, FP, and FN, we compute the following metrics: accuracy, precision, recall, and F1-score (F1) as defined in Equations 20-23.
@@ VI-D Research Questions
Firstly, we analyze RQ1. Figure 6 presents a trend chart of the evaluation metrics as the model is evaluated with $10,000$ smart contracts. The metrics AC, PR, and RE represent accuracy, precision, and recall, respectively. It is observed that after approximately $7,344$ evaluations, these metrics stabilize. Therefore, the metrics evaluated on $10,000$ smart contracts are deemed reasonable and reliable. Finally, SmartIntentNN achieves an F1-score of $0.8633$, accuracy of $0.9647$, precision of $0.8873$, and recall of $0.8406$, corresponding to the first row of Table II, which represents the best-performing variant of SmartIntentNN.
@@ IX Conclusion
We have proposed SmartIntentNN, the first approach for detecting development intent in smart contracts. The model comprises a pre-trained USE model, an intent highlight model based on K-means clustering, and a DNN integrated with a BiLSTM layer. We trained and evaluated our model on a dataset of over $40,000$ smart contracts, which were cleaned and labeled into ten distinct categories of intents. Experimental results demonstrate that our model surpasses all baselines, achieving an accuracy of $0.9647$, a precision of $0.8873$, a recall of $0.8406$, and an F1-score of $0.8633$. Our work fills a critical gap in smart contract security study by addressing developer intent detection.

### v3

@@ V-C Multi-label Classification
Formulas 14 to 19 illustrate the computation process after a row vector $\bm{f_{i}}$ is input into an LSTM cell, where $i\in\mathbb{N}^{+}$ and $i\leq p$. Here, $\Theta$ denotes the cell state vector, and $\widetilde{\Theta}$ signifies the cell input activation vector. The initial state $\Theta_{0}$ is initialized as a zero vector with a dimensionality of $h$ (hidden units per LSTM cell). The various gates within the LSTM cell are represented as follows: $\mathrm{G}^{\mathsf{f}}_{i}$ is the forget gate, $\mathrm{G}^{\mathsf{u}}_{i}$ is the update (or input) gate, and $\mathrm{G}^{\mathsf{o}}_{i}$ is the output gate. The weight matrices $\bm{W}\in\mathbb{R}^{h\times m}$ and $\bm{U}\in\mathbb{R}^{h\times h}$, along with the bias vector $\bm{b}\in\mathbb{R}^{h}$, are parameters learned during the training phase. The activation functions denoted by $\alpha$, $\beta$, and $\gamma$ can be $\mathrm{tanh}$ or $\mathrm{sigmoid}$. The hidden state vector $\bm{h_{i}}$, which is also the output of the LSTM cell, has a dimensionality of $h$, with the initial hidden state $\bm{h_{0}}$ being initialized as a zero vector. The operator $\odot$ denotes the Hadamard product (element-wise product) [49].
@@ VI-B Evaluation Metrics
Using the values of TP, TN, FP, and FN, we compute the following metrics: accuracy, precision, recall, and F1-score (F1) as defined in Equations 21-24.
@@ VI-D Research Questions
Firstly, we analyze RQ1. Fig. 6 presents a trend chart of the evaluation metrics as the model is evaluated with $10,000$ smart contracts. The metrics AC, PR, and RE represent accuracy, precision, and recall, respectively. It is observed that after approximately $7,344$ evaluations, these metrics stabilize. Therefore, the metrics evaluated on $10,000$ smart contracts are deemed reasonable and reliable. Finally, SmartIntentNN achieves an F1-score of $0.8633$, accuracy of $0.9647$, precision of $0.8873$, and recall of $0.8406$, corresponding to the first row of Table II, which represents the best-performing variant of SmartIntentNN.
@@ IX Conclusion
We have proposed SmartIntentNN, the first model for detecting development intent in smart contracts. The model comprises a pre-trained USE model, an intent highlight model based on K-means clustering, and a DNN integrated with a BiLSTM layer. We trained and evaluated our model on a dataset of over $40,000$ smart contracts, which were cleaned and labeled into ten distinct categories of intents. Experimental results demonstrate that our model surpasses all baselines, achieving an accuracy of $0.9647$, a precision of $0.8873$, a recall of $0.8406$, and an F1-score of $0.8633$. Our work fills a critical gap in smart contract security study by addressing developer intent detection.
Acknowledgment
This work was supported by the Macao Science and Technology Development Fund under Grant 0161/2023/RIA.